OpenStudy (anonymous):

what is thee 32nd term of the arithmetic sequence where a1=12 and a13= -60 A) -192 B) -186 C) -180 D) -174

OpenStudy (anonymous):

To find 32nd term first find d using formula an = a1 + (n-1)d Plug an = 13 a1 = 12 & n = 13 & find d , the again use this formula for a32 = 12 + (32 - 1)*d